# Finite and full scale localization for the multi-frequency quasi-periodic CMV matrices

### v2

@@ 3.5 Finite Scale Localization
for any $m\in[0,n-l]\backslash I$. Then for any $(\omega,x)\in\mathbb{T}^{d}(p,q)\times\mathbb{T}^{d}$ satisfying $|x-x_{0}|,|\omega-\omega_{0}|<\exp(-l)$ and any eigenvalue $|z_{j}^{[0,n-1]}(\omega,x)-z_{0}|<\exp(-l)$, we have
@@
Take $x,\omega,z=z_{j}^{[0,n-1]}(\omega,x)$ satisfying the assumptions and $s\in[0,n-1]\backslash I$ such that $d=\mathrm{dist}(s,I)\geq l^{2/\nu}$. Suppose $k\in[0,n^{\prime}]$. Let $J=[s-d,s+d]\cap[0,n-1]=:[t,n^{\prime}]$. Obviously, $|J|\geq l^{2/\nu}$.
@@
Let $\tilde{I}=\{s\in[0,n-1]:\mathrm{dist}(s,I)<l^{2/\nu}\}$. According to Lemma 3.27, one can obtain
@@ 4 Removing Double Resonances by Using Semialgebraic Sets
Let ~$\tilde{\mathcal{E}}_{[0,n-1]}^{\beta,\eta}$ be the matrix with this truncated Verblunsky coefficient and the refined boundary condition $\beta,\eta\in\partial\mathbb{D}$. If we let $z_{j}^{[0,n-1]}(x,\omega)$ and $\tilde{z}_{j}^{[0,n-1]}(x,\omega)$ be the eigenvalues of $\mathcal{E}_{[0,n-1]}^{\beta,\eta}$ and ~$\tilde{\mathcal{E}}_{[0,n-1]}^{\beta,\eta}$ respectively, then
@@
$\frac{1}{nJ}\sum_{j=1}^{J}\log\|\tilde{M}_{n}(\omega,z;x+jw)\|_{HS}\geq\frac{\gamma}{2}$ (4.5)
@@
$\log|\varphi_{[0,n-1]}^{\beta,\eta}(\omega,z;x)|\leq\frac{1}{J}\sum_{j=1}^{J}\log\|\tilde{M}_{n}(\omega,z;x+jw)\|_{HS}-n^{1-\tau/2}/2.$ (4.6)
@@
for all $n\in\mathop{\cup}\limits_{j=1}^{s-1}\mathcal{N}_{k(s-1)+j}$. Taking $i=k(s-1)+j$, the conclusion follows. ∎
@@
Let $\nu$ be as in LDT. Assume $x_{0}\in\mathbb{T}^{d}$, $\omega_{0}\in\mathbb{T}^{d}(p,q)$, $z_{0}\in\partial\mathbb{D}$ and $L(\omega_{0},z_{0})>\gamma>0$. Let $l,n$ be integers such that $(2\log n)^{1/\nu}\leq Cl\leq n$. Then for any $n\geq N_{0}(p,q,\gamma)$ there exists a set $\mathcal{B}_{n,\omega_{0},z_{0}}$, $\mathrm{mes}(\mathcal{B}_{n,\omega_{0},z_{0}})<\exp(-Cl^{\nu}/2)$ such that for any $x\in\mathbb{T}^{d}\backslash\mathcal{B}_{n,\omega_{0},z_{0}}$ and any $(\omega,z)\in\mathbb{T}^{d}\times\partial\mathbb{D}$, $|x-x_{0}|$, $|\omega-\omega_{0}|$, $|z-z_{0}|<\exp(-l)$, we have
@@
$\frac{1}{|\Lambda_{1}|J}\sum_{j=1}^{J}\log\|\tilde{M}_{|\Lambda_{1}|}(\omega,z;x+j\omega)\|_{HS}\geq\frac{\gamma}{2},$
(~$\|(\tilde{\mathcal{E}}_{\Lambda_{0}}^{\beta,\eta}(\omega,x_{0})-z)^{-1}\|>\frac{1}{2}\exp(\bar{l}^{\varpi}),\,\,\mathrm{and}\,\,\|(\tilde{\mathcal{E}}_{\Lambda_{1}}^{\beta,\eta}(\omega,x)-z)^{-1}\|>\frac{1}{2}\exp(\bar{l}^{\varpi}).$
@@
Note that $\{\omega:(\omega,\{x_{0}+t\omega\})\in\mathcal{S}_{1}\}\subset\mathrm{Proj}_{\omega}\mathcal{S}_{1}$. From the assumptions, there are less than $C\bar{l}^{4}$ possible choices for $\Lambda_{0},\Lambda_{1}$. Then we only need to estimate the measure of $\Omega_{\bar{l},t_{0},x_{0}}$. Note that the set of $(\omega,\{x_{0}+t\omega\})$, $\omega\in[0,1]^{d}$, is contained in a union of planes $L_{i}$, $i\leq|t|^{d}$. The planes $L_{i}$ are parallel to $(\omega,t\omega\})$, $\omega\in\mathbb{R}^{d}$, then
@@
We use the notation and assumptions of Lemma 4.12. For any $\bar{l}\geq\bar{l}_{0}(\alpha_{i},p,q,\varpi,\gamma)$ and $1\leq t_{0}\leq\exp(c_{0}\bar{l}^{\varpi\nu})$, there exists a set $\Omega_{\bar{l},t_{0}}$, $\mathrm{mes}(\Omega_{\bar{l},t_{0}})<\bar{l}^{C(p,q)}t_{0}^{-1/2}$, such that for any $\omega\notin\Omega_{\bar{l},t_{0}}$ there exists a set $\mathcal{B}_{\bar{l},t_{0},\omega}$, $\mathrm{mes}(\mathcal{B}_{\bar{l},t_{0},\omega})<\bar{l}^{C(p,q)}t_{0}^{-1/2}$, and the following statement holds. For any $\omega\in\mathbb{T}^{d}(p,q)\backslash\Omega_{\bar{l},t_{0}}$, $x\in\mathbb{T}^{d}\backslash\mathcal{B}_{\bar{l},t_{0},\omega}$, $z\in\partial\mathbb{D}$,such that $L(\omega,z)>\gamma$, we have
@@ 5 Removing Double Resonances under NDR Condition
(c) for any $\underline{\omega}\in\mathcal{P}$, $P(\cdot,\underline{\omega})$ has no zeros in \$\mathbb{C}\backslash\mathcal{D}(z_{0},r)$.
@@
Then there exists $r_{0}/2<r<r_{0}$, which depends only on $\mathcal{E}_{0}$, such that $\mathcal{E}_{1}$ and $\mathcal{E}_{0}$ have the same number of eigenvalues in the disk $\mathcal{D}(z_{0},r)\cap\partial\mathbb{D}$. Moreover, neither $\mathcal{E}_{1}$ nor $\mathcal{E}_{0}$ have eigenvalues in the region
@@
(c) for any $(\omega,x)\in\mathcal{P}$, $P(\omega,x,\cdot)$ has no zeros in \$\mathbb{C}\backslash\mathcal{D}(z_{0},r)$;
@@
and for any $(\omega,x)\in(\mathbb{T}^{d}(p,q)\times\mathbb{T}^{d})\backslash\mathcal{B}_{\bar{l},t}$, $z\in\partial\mathbb{D}$,
@@
According to Lemma 3.7, $L(\omega_{0},z)>\frac{\gamma}{2}$ for any $|z-z_{0}|<\exp(-l)$ (provided $l$ is large enough). Let $\Omega_{\bar{l},t_{0}}$ $\mathcal{B}_{\bar{l},t_{0},\omega_{0}}$ be the sets from Corollary 4.13, then
@@
Let $\tau,\nu$ be as in LDT, $1\leq l\leq\bar{l}\leq\exp(l)$, $1\leq K\leq\bar{l}^{\frac{1-\tau}{10}}$, and $\exp(l^{2})\leq n$. Let $\Lambda_{j}$, $j=0,1$, be intervals in $\mathbb{Z}$ such that $\bar{l}/10\leq|\Lambda_{j}|\leq 10\bar{l}$, $0\in\Lambda_{j}$, $j=0,1$. For $l\geq l_{0}(\alpha_{i},p,q,\gamma)$ there exists $\varrho=\varrho(p,q)$ such that if $n\leq\exp(\bar{l}^{\varrho})$, there exists a set $\Omega_{n,\bar{l}}$, $\mathrm{mes}(\Omega_{n,\bar{l}})<2\exp(-l^{2}/4)$, and for every $\omega\in\mathbb{T}^{d}(p,q)\backslash\Omega_{n,\bar{l}}$ there exists a set $\mathcal{B}_{n,\bar{l},\omega}$, $\mathrm{mes}(\mathcal{B}_{n,\bar{l},\omega})<\exp(-\bar{l}^{\varrho})$ such that the following statement holds. For any $\omega\in\mathbb{T}^{d}(p,q)\backslash\Omega_{n,\bar{l}}$, $x\in\mathbb{T}^{d}\backslash\mathcal{B}_{n,\bar{l},\omega}$, $z\in\partial\mathbb{D}$, and $m_{0},m_{1}\in[0,n-1]$ such that $|m_{0}-m_{1}|\geq\exp(l^{2})$, we have that if $L(\omega,z)>\gamma$ and $\Lambda_{j}$, $j=0,1$, are $(K,l,\frac{1}{2})$-NDR intervals with respect to $x+m_{j}\omega$, $\omega$, $z$, then
@@ 6 Full Scale Localization
(b) If in addition there exists $\hat{\varepsilon}>\varepsilon$ such that the subspace of the eigenvectors of $A$ with eigenvalues falling into the interval $(E-\hat{\varepsilon},E+\hat{\varepsilon})$ is at most of dimension one, then there exists a normalized eigenvector $\psi$ of $A$ with an eigenvalue $E_{0}\in(E-\varepsilon,E+\varepsilon)$, such that
@@
and the following holds for any $\omega\in\mathbb{T}^{d}(p,q)\backslash\Omega_{n}$, $x\in\mathbb{T}^{d}(p,q)\backslash\mathcal{B}_{n,\omega}$ and $z\in\partial\mathbb{D}$ such that $L(\omega,z)>\gamma$, where $\varrho$ is as in Lemma 5.10. Let $1\ll\underline{C}_{1}\ll\overline{C}_{1}\ll\cdots\ll\underline{C}_{(s-1)^{2}}\ll\overline{C}_{(s-1)^{2}}$, $s=2^{2d+1}$, be constants such that the sets
@@
(there exists $k\in\{1,\ldots,(s-1)^{2}\}$ such that the above inequality holds for all $m^{\prime}\in\mathcal{N}_{k}$) or there exists $k=k(m_{0}^{\prime},m^{\prime})$ such that $[-\bar{l}_{k},\bar{l}_{k}]$ is $(K_{k},l,\frac{1}{2})$-NDR with respect to both $x+(m_{0}^{\prime}-1)\omega,\omega,z$ and $x+(m^{\prime}-1)\omega,\omega,z$.
@@
where $\Lambda(m^{\prime})$ is either $[0,l-1]$ or $[-\bar{l}_{k},\bar{l}_{k}]$ with $k=k(m_{0}^{\prime},m^{\prime})$ as above.
According to the definition of NDR intervals, if $[-\bar{l}_{k},\bar{l}_{k}]$ is NDR with respect to $x+(m_{0}^{\prime}-1)\omega,\omega,z$, then $[-\bar{l}_{k},\bar{l}_{k}]\cap([0,n-1]-m_{0}^{\prime}+1)$ is also NDR with respect to $x+(m_{0}^{\prime}-1)\omega,\omega,z$. Therefore, Lemma 5.10 guarantees that if $\Lambda(m^{\prime})=[-\bar{l}_{k},\bar{l}_{k}]$, then (6.2) also holds for any subintervals of $[-\bar{l}_{k},\bar{l}_{k}]$ with length exceeding than $\bar{l}_{k}/10$.
Consider $m\in[0,n-\bar{l}]$ such that $|m-m_{0}|\geq\bar{l}+2\exp{l^{2}}$. Then $[m,m+\bar{l}-1]\subset\{m^{\prime}:|m^{\prime}-m_{0}^{\prime}|\geq\exp(l^{2})\}$. Every point in $[m,m+\bar{l}-1]$ is contained in an interval of the form $\Lambda(m^{\prime})\cap[m,m+\bar{l}-1]$ for some $m^{\prime}\in[m-1,m+\bar{l}-1-l]$, on which the large deviations estimate applies. Note that we aimed to ensure that $m^{\prime}+[1,l]\subset[m,m+\bar{l}-1]$, since if a large deviations estimate holds on $m^{\prime}+[1,l]$, it does not necessarily hold on $m^{\prime}+[1,l]\subset[m,m+\bar{l}-1]$. On the other hand, we already noted that when a large deviations estimate holds on $m^{\prime}+[-\bar{l}_{k},\bar{l}_{k}]$ it also holds on $(m^{\prime}+[-\bar{l}_{k},\bar{l}_{k}])\cap[m,m+\bar{l}-1]$. Since $\bar{l}\geq(\exp(\overline{C}_{(s-1)^{2}}(\log n)^{3\varepsilon\nu/2}))^{2/\nu}$, then
@@
Let $\varepsilon\in(0,\frac{1}{5})$, $\nu$ as in LDT, and $\Omega_{n}$, $\mathcal{B}_{n,\omega}$ as in Lemma 9.1. For any $n\geq N_{0}(\alpha_{i},p,q,\gamma,\varepsilon)$, $\omega_{0}\in\mathbb{T}^{d}(p,q)\backslash\Omega_{n}$, $x_{0}\in\mathbb{T}^{d}\backslash\mathcal{B}_{n,\omega}$, and any eigenvalue $z_{0}=z_{j}^{[0,n-1]}(\omega_{0},x_{0})$, such that $L(\omega_{0},z_{0})>\gamma$, there exists an interval $I=I(\omega_{0},z_{0},x_{0})\subset[0,n-1]$,
@@
Then $m_{0}$ satisfies the assumptions of Lemma 6.3 because otherwise we can derive a contradiction based on the proof of Lemma 3.23. Let $\bar{l}=\lceil(\exp(\overline{C}_{(s-1)^{2}}(\log n)^{3\varepsilon\nu/2}))^{2/\nu}\rceil$. According to Lemma 6.3 (with $s$ instead of $m$), (6.1) holds for $s\in[0,n-1]\backslash I$, $I=[0,n-1]\cap[m_{0}-\bar{l}-2\exp(l^{2}),m_{0}+\bar{l}+2\exp(l^{2})]$, $l=\lceil(\log n)^{2\varepsilon}\rceil$. Then we can obtain the conclusion by applying Theorem 3.27 (with $\bar{l}$ instead of $l$). ∎
@@
Let $\varepsilon\in(0,\frac{1}{5})$ and $\Omega_{n}$, $\mathcal{B}_{n,\omega}$ be as in Lemma 6.4. For any $n\geq N_{0}(\alpha_{i},p,q,\gamma,\varepsilon)$, $\omega_{0}\in\mathbb{T}^{d}(p,q)\backslash\Omega_{n}$, $x_{0}\in\mathbb{T}^{d}(p,q)\backslash\mathcal{B}_{n,\omega}$, and any eigenvalue $z_{0}=z_{j}^{[0,n-1](\omega_{0},x_{0})}$, such that $L(\omega_{0},z_{0})>\gamma$, if we take $I=I(\omega_{0},z_{0},x_{0})$ as in Lemma 6.4, then
@@
According to the proof of Lemma 6.4, (6.1) holds for $s\in[0,n-1]\backslash I$, $I=[0,n-1]\cap[m_{0}-\bar{l}-2\exp(l^{2}),m_{0}+\bar{l}+2\exp(l^{2})]$, $l=\lceil(\log n)^{2\varepsilon}\rceil$. Then the conclusion follows from Lemma 3.28. ∎
In what follows, when we deal with the difference between two vectors of different dimensions, we make the dimension of a low dimensional vector equal to the dimension of a high dimensional vector by adding the 00 element. For example, $\xi=(x_{-5},x_{-4},\ldots,x_{-1},x_{0},x_{1},\ldots,x_{4},x_{5})$, $\upsilon=(y_{-7},y_{-6},\ldots,y_{-1},y_{0},y_{1},\ldots,y_{6},y_{7}),$ $n>m$, $\upsilon-\xi=(y_{-7},y_{-6},y_{-5}-x_{-5},\ldots,y_{-1}-x_{-1},y_{0}-x_{0},y_{1}-x_{1},\ldots,y_{5}-x_{5},y_{6},y_{7})$. When the columns of a matrix are greater than the rows of a vector, we deal with their product in the same way.
@@
Let $\varepsilon\in(0,\frac{1}{5})$. With $\Omega_{n}$, $\mathcal{B}_{n,\omega}$ as in Lemma 6.4, applied on $[-(n-1),n-1]$ instead of $[0,n-1]$, let $\hat{\Omega}_{N_{0}}=\mathop{\cup}\limits_{n\geq N_{0}}\Omega_{n}$, ^$\hat{\mathcal{B}}_{N_{0},\omega}=\mathop{\cup}\limits_{n\geq N_{0}}\mathcal{B}_{n,\omega}$. For $n\geq N_{0}(\alpha_{i},p,q,\gamma,\varepsilon)$, $\omega\in\mathbb{T}^{d}(p,q)\backslash\hat{\Omega}_{N_{0}}$, $x\in\mathbb{T}^{d}\backslash\hat{\mathcal{B}}_{N_{0},\omega}$, if $L(\omega,z_{j}^{[-(n-1),n-1]}(\omega,x))>\frac{3\gamma}{2}$ and $I=I(\omega,z_{j}^{[-(n-1),n-1]}(\omega,x),x)$ then for any $n\leq n^{\prime}\leq\exp((\log n)^{\frac{1}{10\varepsilon}})$ there exists $j_{n^{\prime}}$, such that

### v3

@@ 3.5 Finite Scale Localization
for any m[$\Im\in[0,n-l]\backslash$ I. Then for any $(\omega,x)\in\mathbb{T}^{d}(p,q)\times\mathbb{T}^{d}$ satisfying $|x-x_{0}|,|\omega-\omega_{0}|<\exp(-l)$ and any eigenvalue $|z_{j}^{[0,n-1]}(\omega,x)-z_{0}|<\exp(-l)$, we have
@@
Take $x,\omega,z=z_{j}^{[0,n-1]}(\omega,x)$ satisfying the assumptions and s[$\Is\in[0,n-1]\backslash$ I such that $d=\mathrm{dist}(s,I)\geq l^{2/\nu}$. Suppose $k\in[0,n^{\prime}]$. Let $J=[s-d,s+d]\cap[0,n-1]=:[t,n^{\prime}]$. Obviously, $|J|\geq l^{2/\nu}$.
@@
To complement the finite scale localization of eigenfunctions, we next establish a separation result for eigenvalues.
@@
Let $\tilde{I}=\{s\in[0,n-1]:\mathrm{dist}(s,I)<l^{2/\nu}\}$. According to Theorem 3.26, one can obtain
@@ 4 Removing Double Resonances by Using Semialgebraic Sets
Let $\tilde{\mathcal{E}}_{[0,n-1]}^{\beta,\eta}$ be the matrix with this truncated Verblunsky coefficient and the refined boundary condition $\beta,\eta\in\partial\mathbb{D}$. If we let $z_{j}^{[0,n-1]}(x,\omega)$ and $\tilde{z}_{j}^{[0,n-1]}(x,\omega)$ be the eigenvalues of $\mathcal{E}_{[0,n-1]}^{\beta,\eta}$ and $\tilde{\mathcal{E}}_{[0,n-1]}^{\beta,\eta}$ respectively, then
@@
We first recall the definition of semialgebraic sets, which are finite unions of sets defined by polynomial equalities and inequalities. This structure is critical for measuring exceptional resonant sets.
@@
To control the number of resonant points along orbits of the shift map, we use the following lemma, which bounds the density of semialgebraic sets under Diophantine frequency shifts.
@@
Next, we extend this density control to intersections of semialgebraic sets, which is necessary for handling multiple resonant conditions simultaneously.
@@
To control resonant frequencies, we use the following lemma, which bounds the measure of frequencies for which orbits pass through a semialgebraic set. This is key for excluding exceptional frequencies that support double resonances.
@@
We now show that the set of phases where the transfer matrix norm deviates from its LDT average is contained in a semialgebraic set of small measure.
@@
We extend the LDT average to time-averages along the shift orbit, ensuring that the average of the transfer matrix norm over a long orbit is close to the finite scale Lyapunov exponent.
@@
We characterize the set of resonant $(\omega,z,x)$ where the characteristic determinant is small as a semialgebraic set of small measure.
@@
$\frac{1}{nJ}\sum_{j=1}^{J}\log\|\tilde{M}_{n}(\omega,z;x+jw)\|_{\mathrm{HS}}\geq\frac{\gamma}{2}$ (4.5)
@@
$\log|\varphi_{[0,n-1]}^{\beta,\eta}(\omega,z;x)|\leq\frac{1}{J}\sum_{j=1}^{J}\log\|\tilde{M}_{n}(\omega,z;x+jw)\|_{\mathrm{HS}}-n^{1-\tau/2}/2.$ (4.6)
@@
We use the semialgebraic characterization to show that for most frequencies, resonant phases are rare along orbits.
@@
We extend the previous result to two resonant phases, ensuring that for most frequencies, two distinct resonant phases cannot have long orbits without non-resonant points.
@@
for all $n\in\mathop{\cup}\limits_{j=1}^{s-1}\mathcal{N}_{k(s-1)+j}$. Taking i=k(s1)+ji=k(s-1)+j, the conclusion follows. ∎
@@
[…]
@@
We extend the Wegner estimate to control resonances between two distant intervals, showing that for most frequencies and phases, at least one of the intervals has a bounded Green’s function (i.e., no resonance). This directly eliminates double resonances.
@@
$\frac{1}{|\Lambda_{1}|J}\sum_{j=1}^{J}\log\|\tilde{M}_{|\Lambda_{1}|}(\omega,z;x+j\omega)\|_{\mathrm{HS}}\geq\frac{\gamma}{2},$
$\|(\tilde{\mathcal{E}}_{\Lambda_{0}}^{\beta,\eta}(\omega,x_{0})-z)^{-1}\|>\frac{1}{2}\exp(\bar{l}^{\varpi}),\,\,\mathrm{and}\,\,\|(\tilde{\mathcal{E}}_{\Lambda_{1}}^{\beta,\eta}(\omega,x)-z)^{-1}\|>\frac{1}{2}\exp(\bar{l}^{\varpi}).$
@@
Note that $\{\omega:(\omega,\{x_{0}+t\omega\})\in\mathcal{S}_{1}\}\subset\mathrm{Proj}_{\omega}\mathcal{S}_{1}$. From the assumptions, there are less than $C\bar{l}^{4}$ possible choices for $\Lambda_{0},\Lambda_{1}$. Then we only need to estimate the measure of $\Omega_{\bar{l},t_{0},x_{0}}$. Note that the set of $(\omega,\{x_{0}+t\omega\})$, $\omega\in[0,1]^{d}$, is contained in a union of planes $L_{i}$, $i\leq|t|^{d}$. The planes $L_{i}$ are parallel to $(\omega,t\omega)$, $\omega\in\mathbb{R}^{d}$, then
@@
We further extend the result to hold for most phases, ensuring that for most frequencies and phases, double resonances between distant intervals are excluded.
@@
[…]
@@ 5 Removing Double Resonances under NDR Condition
(c) for any $\underline{\omega}\in\mathcal{P}$, $P(\cdot,\underline{\omega})$ has no zeros in $\mathbb{C}\backslash\mathcal{D}(z_{0},r)$.
We also use matrix perturbation theory to ensure that eigenvalues of NDR intervals are stable under small perturbations of frequency and phase.
@@
Then there exists $r_{0}/2<r<r_{0}$, which depends only on $\mathcal{E}_{0}$, such that $\mathcal{E}_{1}$ and $\mathcal{E}_{0}$ have the same number of eigenvalues in the arc $\mathcal{D}(z_{0},r)\cap\partial\mathbb{D}$. Moreover, neither $\mathcal{E}_{1}$ nor $\mathcal{E}_{0}$ have eigenvalues in the region
@@
Using the unitary perturbation result (Corollary 5.4), we show that eigenvalues of NDR intervals are isolated from a fixed circle under small perturbations.
@@
We also use matrix perturbation theory to ensure that eigenvalues of NDR intervals are stable under small perturbations of frequency and phase.
@@
(c) for any $(\omega,x)\in\mathcal{P}$, $P(\omega,x,\cdot)$ has no zeros in $\mathbb{C}\backslash\mathcal{D}(z_{0},r)$;
@@
We use the following property of resultants to link non-vanishing resultants to uniform lower bounds on the polynomials, ensuring no common zeros.
@@
Finally, we combine the NDR factorization, resultant bounds, and semialgebraic set tools to eliminate double resonances between two distant NDR intervals. We show that for most frequencies and phases, at least one of the intervals has a non-vanishing polynomial factor (no resonance), ruling out common eigenvalues.
@@
[…]
@@
According to Lemma 3.7, $L(\omega_{0},z)>\frac{\gamma}{2}$ for any $|z-z_{0}|<\exp(-l)$ (provided $l$ is large enough). Let $\Omega_{\bar{l},t_{0}}$, $\mathcal{B}_{\bar{l},t_{0},\omega_{0}}$ be the sets from Corollary 4.13, then
@@
We extend the above result to all pairs of distant points in a large interval, ensuring that double resonances are excluded uniformly across the interval. This uniform exclusion is critical for extending finite scale localization to full scale localization.
@@
[…]
@@ 6 Full Scale Localization
[…]
@@
Next, we show that for most frequencies and phases, eigenfunctions of large finite intervals are localized to a small subinterval, with exponential decay away from this localization center.
@@
[…]
@@
(there exists $k\in\{1,\ldots,(s-1)^{2}\}$ such that the above inequality holds for all $m^{\prime}\in\mathcal{N}_{k}$) or there exists $k=k(m_{0}^{\prime},m^{\prime})$ such that $\bar{l}_{k},\bar{l}_{k}]$ is $(K_{k},l,\frac{1}{2})$-NDR with respect to both $x+(m_{0}^{\prime}-1)\omega,\omega,z$ and $x+(m^{\prime}-1)\omega,\omega,z$.
@@
where $\Lambda(m^{\prime})$ is either [0,l1][0,l-1] or $\bar{l}_{k},\bar{l}_{k}]$ with $k=k(m_{0}^{\prime},m^{\prime})$ as above.
According to the definition of NDR intervals, if $\bar{l}_{k},\bar{l}_{k}]$ is NDR with respect to $x+(m_{0}^{\prime}-1)\omega,\omega,z$, then $\bar{l}_{k},\bar{l}_{k}]\cap([0,n-1]-m_{0}^{\prime}+1)$ is also NDR with respect to $x+(m_{0}^{\prime}-1)\omega,\omega,z$. Therefore, Lemma 5.10 guarantees that if $\Lambda(m^{\prime})=[-\bar{l}_{k},\bar{l}_{k}]$, then (6.2) also holds for any subintervals of $\bar{l}_{k},\bar{l}_{k}]$ with length exceeding than $\bar{l}_{k}/10$.
Consider $m\in[0,n-\bar{l}]$ such that $|m-m_{0}|\geq\bar{l}+2\exp{l^{2}}$. Then $m,m+\bar{l}-1]\subset\{m^{\prime}:|m^{\prime}-m_{0}^{\prime}|\geq\exp(l^{2})\}$. Every point in $m,m+\bar{l}-1]$ is contained in an interval of the form $\Lambda(m^{\prime})\cap[m,m+\bar{l}-1]$ for some $m^{\prime}\in[m-1,m+\bar{l}-1-l]$, on which the large deviations estimate applies. Note that we aimed to ensure that $m^{\prime}+[1,l]\subset[m,m+\bar{l}-1]$, since if a large deviations estimate holds on $m^{\prime}+[1,l]$, it does not necessarily hold on $m^{\prime}+[1,l]\cap[m,m+\bar{l}-1]$. On the other hand, we already noted that when a large deviations estimate holds on $m^{\prime}+[-\bar{l}_{k},\bar{l}_{k}]$ it also holds on $(m^{\prime}+[-\bar{l}_{k},\bar{l}_{k}])\cap[m,m+\bar{l}-1]$. Since $\bar{l}\geq(\exp(\overline{C}_{(s-1)^{2}}(\log n)^{3\varepsilon\nu/2}))^{2/\nu}$, then
@@
Building on the above lemma, we prove that localized eigenfunctions of large finite intervals decay exponentially away from their localization center, with the decay rate independent of the interval size.
@@
[…]
@@
Then $m_{0}$ satisfies the assumptions of Lemma 6.3 because otherwise we can derive a contradiction based on the proof of Lemma 3.23. Let $\bar{l}=\lceil(\exp(\overline{C}_{(s-1)^{2}}(\log n)^{3\varepsilon\nu/2}))^{2/\nu}\rceil$. According to Lemma 6.3 (with $s$ instead of $m$), (6.1) holds for s[$\Is\in[0,n-1]\backslash$ I, $I=[0,n-1]\cap[m_{0}-\bar{l}-2\exp(l^{2}),m_{0}+\bar{l}+2\exp(l^{2})]$, $l=\lceil(\log n)^{2\varepsilon}\rceil$. Then we can obtain the conclusion by applying Theorem 3.26 (with $\bar{l}$ instead of $l$). ∎
To ensure eigenpair convergence, we show that eigenvalues corresponding to localized eigenfunctions are well-separated. This prevents eigenvalue crossings as the interval expands, ensuring each localized eigenfunction converges to a unique eigenfunction of the infinite matrix.
@@
[…]
@@
According to the proof of Lemma 6.4, (6.1) holds for s[$\Is\in[0,n-1]\backslash$ I, $I=[0,n-1]\cap[m_{0}-\bar{l}-2\exp(l^{2}),m_{0}+\bar{l}+2\exp(l^{2})]$, $l=\lceil(\log n)^{2\varepsilon}\rceil$. Then the conclusion follows from Lemma 3.27. ∎
In what follows, when we deal with the difference between two vectors of different dimensions, we make the dimension of a low dimensional vector equal to the dimension of a high dimensional vector by adding the 0 element. For example, $\xi=(x_{-5},x_{-4},\ldots,x_{-1},x_{0},x_{1},\ldots,x_{4},x_{5})$, $\upsilon=(y_{-7},y_{-6},\ldots,y_{-1},y_{0},y_{1},\ldots,y_{6},y_{7}),$ $n>m$, $\upsilon-\xi=(y_{-7},y_{-6},y_{-5}-x_{-5},\ldots,y_{-1}-x_{-1},y_{0}-x_{0},y_{1}-x_{1},\ldots,y_{5}-x_{5},y_{6},y_{7})$. When the columns of a matrix are greater than the rows of a vector, we deal with their product in the same way.
@@
Let $\varepsilon\in(0,\frac{1}{5})$. With $\Omega_{n}$, $\mathcal{B}_{n,\omega}$ as in Lemma 6.4, applied on [(n1),n1][-(n-1),n-1] instead of [0,n1][0,n-1], let $\hat{\Omega}_{N_{0}}=\mathop{\cup}\limits_{n\geq N_{0}}\Omega_{n}$, $\hat{\mathcal{B}}_{N_{0},\omega}=\mathop{\cup}\limits_{n\geq N_{0}}\mathcal{B}_{n,\omega}$. For $n\geq N_{0}(\alpha_{i},p,q,\gamma,\varepsilon)$, $\omega\in\mathbb{T}^{d}(p,q)\backslash\hat{\Omega}_{N_{0}}$, $x\in\mathbb{T}^{d}\backslash\hat{\mathcal{B}}_{N_{0},\omega}$, if $L(\omega,z_{j}^{[-(n-1),n-1]}(\omega,x))>\frac{3\gamma}{2}$ and $I=I(\omega,z_{j}^{[-(n-1),n-1]}(\omega,x),x)$ then for any $n\leq n^{\prime}\leq\exp((\log n)^{\frac{1}{10\varepsilon}})$ there exists $j_{n^{\prime}}$, such that